Calculate the entropy of mixing when 2.00 mol of H2 are mixed with 3.00 mol of N2 assuming that no chemical reaction accur between them
1
Expert's answer
2020-09-22T03:34:55-0400

Ans.

Total mole number of hydrogen and nitrogen = (2+3)= 5 mol

Mole fraction of H2 = 2/5

Mole fraction of N2 = 3/5

Entropy of mixing = -5R{(2/5)ln(2/5) + (3/5)ln(3/5)} = 3.36R = 27.94J/mol·K